We took a look at the Xend green, and we weren't able to make it lock. We improved the alignment a little bit, and when we looked at the error signal, it looked nice and PDH-y, but for whatever reason, the cavity won't catch lock.
While aligning the green to the arm, Jamie noticed that the reflection from the intracavity power (not the prompt reflection) was not overlapping with the input beam or prompt reflection. This means that the cavity axis and the input green beam were not co-linear. I adjusted the BS and ITMX to get the IR transmitted beam (which had been near clipping on the top edge of the first (2 inch) optic it sees out of the vacuum) back near the input green beam spot on the combining beam splitter. Then we continued tweaking the green alignment until we saw nice TEM00 flashes in the cavity. The SNR of the error signal increased significantly after this work, since the cavity buildup was much higher. But alas, still no lock.